# UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality

### v1

@@ 4.2 Experiment Design
Consistency. We conduct a case study on the “harmless” vs. “humor” task, appending the requirement “<max_utility_idx> i” to each prompt to encourage the model to follow the first preference (i.e., maximizing 9th utility function). We then use reward models to evaluate the generated responses. We provide a detailed analysis in Appendix C.3 to present the distribution of normalized utility values for each utility function over our evaluation data. Notably, the 9th utility function achieves the highest mean normalized value, indicating that the fine-tuned model aligns its responses with the specified preference requirement. This demonstrates the model’s ability to selectively prioritize and adhere to user-designated objectives.
Number of utility functions. We vary the number of utility functions over the set 5, 10, 15 and compare the resulting empirical Pareto fronts in Figure 4. We leave the results of the remaining four tasks in Appendix C.4. With 5 utility functions, the generated Pareto front is slightly inferior to those produced with 10 and 15 utility functions. Specifically, the front with 5 utility functions captures a narrower range of optimal trade-offs across the objectives. However, when the number of utility functions is increased to 10 or 15, the resulting Pareto fronts exhibit comparable performance. Increasing the number of utility functions beyond 10 does not yield substantial gains. These results demonstrate that our method is insensitive to the precise number of utility functions, provided that a sufficient number (around 10) is used to effectively cover the preference space. In practical terms, employing 10 utility functions strikes a good balance between performance and computational efficiency, as further increases to 15 do not produce notable improvements.
Figure 4: Ablation study of different numbers of utility functions on “humor” vs. “helpful” and “deberta” vs. “faithful” tasks.
Variants of utility functions. We explored the impact of using linear utility functions, designed as $w^{T}x$ where $||w||_{2}=1$, and compared the performance with our original approach across all tasks. The generated Pareto fronts, as shown in Figure 5 and Appendix C.5, illustrate that our method continues to deliver strong results even with linear utility functions. However, there is still a performance gap between the results achieved with distributional Pareto-optimal utility functions and those obtained using linear functions. While linear functions still produce reasonably good trade-offs, they fail to capture the full range of optimal solutions attainable by our method, which leverages distributional utility representations. This highlights the advantage of our approach in handling more complex utility functions, demonstrating its superior flexibility in navigating multi-objective optimization.
Figure 5: Ablation study of using linear utility functions.
Constraint satisfaction and variance objective. In our case study on the “harmless” vs. “humor” task, we use 2,000 inference samples and compute both metrics for our method, Rewarded Soup, and RiC. As summarized in Table 3, our method achieves the highest overall constraint satisfaction ratio and exhibits the largest variance objective value, outperforming both Rewarded Soup and RiC. This indicates that our approach offers a more comprehensive distributional coverage, effectively handling diverse user preferences.
  Method Constraint Satisfaction Variance
RiC 0.5637 0.0662
Rewarded Soups 0.6744 0.1631
Ours 0.7184 0.5160
Table 3: Comparison of Methods for Constraint satisfaction and variance objective.
@@ 5 Conclusion
We propose Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel multi-objective LLM alignment framework. By transforming user-specified preference into symbolic tokens through a diverse ensemble of strictly increasing, non-linear utility functions, UC-MOA circumvents the numerical sensitivity issues seen in methods like RiC while avoiding the high computational costs of approaches such as MORLHF, Rewarded Soups, and MODPO. Our experiments demonstrate that UC-MOA achieves superior Pareto fronts across complex reward dimensions, evidencing robust and scalable alignment with human values.
Limitations
Representation of Human Preferences. Although UC-MOA effectively captures the distributional properties of human feedback, its performance heavily depends on the quality and diversity of the underlying preference data. Real-world human feedback is often noisy or biased, and addressing such discrepancies requires robust preprocessing and annotation methods. Additionally, certain complex preferences might not be adequately represented by the utility functions, potentially limiting the model’s applicability in diverse contexts, such as safety-critical domains like healthcare and autonomous systems.
Computational and Scaling Challenges. The proposed framework improves efficiency compared to existing methods, but it still requires substantial computational resources when scaling to tasks with numerous objectives or high-dimensional reward spaces. The reliance on large-scale GPUs and the requirement for iterative online updates may limit its adoption in settings with constrained computational budgets. Further optimization of the model architecture and training process could help mitigate these challenges, making the framework more accessible and enabling faster experimentation.
Generalizability Across Tasks. While UC-MOA has demonstrated superior alignment performance in text generation and summary tasks, its applicability to other domains, such as multimodal tasks, remains unexplored. Extending the framework to these areas might introduce additional challenges, including the need to handle dynamic environments and integrate feedback in near real-time. Adapting and validating this approach across a broader range of use cases to evaluate its generalizability and robustness is a potential direction for future research.
Ethical Considerations
While multi-objective optimization and reward modeling can enhance the alignment of large language models, our method must be applied responsibly to avoid harmful outcomes. The potential biases or harmfulness in training data or reward models may mislead the final LLM’s behavior. Careful curation of training and evaluation datasets, combined with auditing of model outputs, is essential to mitigate these risks.
Appendix A Details of Datasets
@@
The HH-RLHF dataset, as described in Bai et al. (2022), was collected to support research on aligning language models using human preferences. The dataset comprises approximately 160,000 prompts each paired with two responses generated by a language model. For every prompt, one response is labeled as “chosen” (preferred by human annotators) and the other as “rejected”. Human raters evaluated paired responses based on criteria for helpfulness and harmlessness. For helpfulness, annotators selected the response that best addressed the given task, whereas for harmlessness (red-teaming) they identified the response as more likely to produce harmful content. This dual-annotation design allows for training reward models that capture subtle aspects of human judgment. Note that we follow Yang et al. (2024d) and use “chosen” data to construct our offline dataset.
@@
The Reddit Summary dataset, introduced in Stiennon et al. (2020), is intended for research in abstractive summarization with human feedback. The dataset consists of approximately 14,900 Reddit posts paired with corresponding summaries. Each entry comprises a post and its summary, where the summary is either the original TL;DR provided by the poster or a human-annotated summary aimed at capturing the post’s main ideas.
Appendix B Implementation Details
B.1 Utility Functions
@@ B.2 Prompt Labeling
B.3 Hyperparameter Choices
Table 5 summarizes the hyper-parameter setting for our experiments.
@@ B.3 Hyperparameter Choices
Table 5: Hyper-parameter settings in our experiments.
B.4 Distributional Metrics
We use two novel metrics as in Cai et al. (2023) for assessing distributional preference: Constraint Satisfaction and Variance Objective.
For Constraint Satisfaction, we define $u(p,\pi)$ to be the probability of return random variance $z\sim\mathbf{Z}(\pi)$ that satisfies the constraint $p$, representing how a policy satisfies the given constraint. We randomly generate each constraint defined by $n$ linear constraints represented by $\{\mathbf{z}|\forall i,\mathbf{w}_{i}^{T}\mathbf{z}\geq c_{i}\}$, where $w_{i}$ is uniformly sampled from the surface $\mathbf{w}_{i}\succeq 0,\mathbf{1}^{T}\mathbf{w}_{i}=1$, and $c_{i}$ is sampled from the distribution of $\mathbf{U}(\min_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z},\max_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z})$, where $\mathbf{Z}$ is the set of all return samples of the final policies learned by DPMORL.
For Variance Objective, we define $u(p,\pi)$ to be a weighted sum of expected return $\mathbb{E}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}\mathbf{z}$ and standard deviation of return $\sqrt{\text{Var}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}[\mathbf{z}]}$. We randomly generate $\mathbf{w}_{i}\in\mathbb{R}^{2K}$, where $\mathbf{w}_{i}\succeq 0$ and $\mathbf{1}^{T}\mathbf{w}_{i}=1$. We let $\mathbf{w}_{i}=[\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}]$, where $\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}\in\mathbb{R}^{K}$, and use $\mathbf{w}_{i}^{1}$ and -$\mathbf{w}_{i}^{2}$ as the weight for expected return and standard derivation of return respectively.
For both of the metrics, we define $u(p,\pi)$ to be the satisfaction score of policy $\pi$ under preference $p$. We generate $M=100$ number of preference $p_{1},p_{2},\cdots,p_{M}$. For both of the Constraint Satisfaction and Variance Objective metric, the mean score is defined as
$\frac{1}{M}\sum_{i=1}^{M}\max_{j\in\{1,2,\cdots,N\}}u(p_{i},\pi_{j})$
The mean score in the above equation measures the best performance of satisfying the preference of the user, averaged in $M=100$ preferences. A higher mean score represents a better ability to meet different preferences.
Appendix C Additional Experiment Results
C.1 Experiment I
@@
C.2 Experiment III
We investigate the impact of the online training stage by comparing three variants of our method: (1) pure offline training (w/o online), (2) one online iteration (w/ online iter 1), and (3) two online iterations (w/ online iter 2). As shown in Figure 8, each subplot corresponds to the Pareto fronts for a different task. Across all four tasks, w/o online (blue line) consistently produces the weakest Pareto front, indicating suboptimal performance when balancing multiple objectives. Introducing a single online iteration (w/ online iter 1, green line) noticeably expands the Pareto front, suggesting that even a brief online training phase can significantly improve multi-reward optimization. Finally, adding a second online iteration (w/ online iter 2, red line) further refines these trade-offs, achieving the most outward (and thus superior) Pareto front among the three variants. These results confirm the necessity of the online stage, demonstrating that multiple online iterations can better align the model with the target objectives.
Figure 8: Ablation study of the online stage on the remaining four tasks.
C.3 Experiment IV
Utility index 0 1 2 3 4 5 6 7 8 9
Mean 0.57 0.57 0.56 0.56 0.57 0.57 0.49 0.55 0.68 0.58
Table 6: Mean normalized value for each utility function.
C.4 Experiment V
We investigate how varying the number of utility functions (5, 10, 15) affects our method’s ability to balance multiple objectives. Figure 10 shows the resulting Pareto fronts across four different tasks, where an outward-shifted curve indicates better overall performance. We observe that using only 5 utility functions (blue line) yields a slightly weaker Pareto front compared to the other two settings. In contrast, the results for 10 (green line) and 15 (red line) utility functions are nearly identical, suggesting that increasing beyond 10 utility functions does not lead to a substantial performance gain. Consequently, our method demonstrates robustness to the exact number of utility functions, so long as it is sufficiently large to capture the relevant trade-offs among the objectives.
Figure 10: Ablation study of different numbers of utility functions on remaining four tasks.
C.5 Experiment VI
@@
We compare the performance of our method when using linear utility functions (Ours (linear)) versus our original, distribution-based approach (Ours (param)) in Figure 11. In each subplot, the curves depict how effectively each method covers the trade-offs between two objectives. Although the linear approach maintains a reasonable Pareto front and thus remains effective for balancing multiple rewards, it consistently trails the distribution-based version in all tasks. This shortfall indicates that purely linear utility functions, while simpler to construct, capture a narrower range of potential trade-offs and consequently yield a less comprehensive frontier. In contrast, our distribution-based design is more flexible, enabling better alignment with diverse user preferences and resulting in a noticeably more outward Pareto front.

### v2

@@ 4.2.2 Quantitative Evaluation of Effectiveness
Consistency in Inference Alignment. To evaluate UC-MOA’s ability to adhere to specified preferences during inference, we embed target utility indices (<max_utility_index> i) into prompts and measure whether the generated responses’ reward profiles align with the targeted utility function. For each generated response, we compute a set of utility values using a suite of ten reward models. To account for scale differences across objectives, the computed reward values are first normalized on a per-utility basis. We then quantify consistency by comparing the expected utility and the value distribution of each utility function with the target objective extracted from the prompt. In a case study on the “harmless” vs. “humor” task, we appended the requirement “<max_utility_idx> i” to each prompt to encourage the model to follow the first preference (i.e., maximizing the 9th utility function). We then used reward models to evaluate the generated responses. A detailed analysis in Appendix E.1 presents the distribution of normalized utility values for each utility function over our evaluation data. Notably, the 9th utility function achieves the highest mean normalized value, indicating that the fine-tuned model aligns its responses with the specified preference requirement. This demonstrates the model’s ability to selectively prioritize and adhere to user-designated objectives.
Pareto Front Analysis. We compare the empirical Pareto fronts generated by each method based on multi-dimensional average test rewards. A method achieving an outer curve demonstrates superior effectiveness in capturing diverse, high-quality trade-offs between objectives. We plot these fronts for all methods across all subtasks. Figure 2 shows the results for the “harmless” vs. “humor” and “deberta” vs. “faithful” tasks (additional results in Appendix E.2). We observe that our method consistently produces an outer Pareto curve that dominates those of MORLHF, Rewarded Soups, and RiC, potentially reflecting its ability to handle complex, non-linear preferences via learned utility functions.
Figure 2: Pareto fronts of “harmless” vs “humor” and “deberta” vs. “faithful” tasks. Each marker on the plotted curves represents the average test rewards corresponding to a particular user preference. Note that a curve that covers a larger area indicates a better performance of the corresponding method.
4.3 Efficiency Analysis
Computational Cost. We measure and compare the training time (GPU hours on 80GB A100) required by each method to achieve its reported performance, providing insight into practical feasibility. Table 1 reports the computational cost. As expected, methods requiring multiple model training or extensive tuning, like Rewarded Soups and MORLHF, exhibit significantly higher costs. RiC is more efficient than these but still requires considerably more time than our approach. UC-MOA, even with two online iterations, demonstrates superior computational efficiency compared to all other multi-objective baselines (MORLHF, Rewarded Soups, and RiC), offering a compelling balance between strong performance and resource usage.
Table 1: Comparison of GPU hours for different methods on Helpful Assistant and Reddit Summary tasks, where the number of preference $M=10$, number of rewards $N=2$.
  Method Helpful Assistant Reddit Summary
  MORLHF 726.5 216.9
Rewarded Soups 69.5 15.9
RiC 20.3 5.7
Ours w/o online 8.9 2.0
Ours w/ online iter1 10.1 2.4
Ours w/ online iter2 11.3 2.8
We also conduct an ablation study to investigate how other factors (i.e., the necessity of the online stage, number of utility functions, and variants of utility functions) influence the performance of our method. The results are presented in Appendix E.3.
5 Limitations
We acknowledge several limitations in the proposed UC-MOA framework. First, the overall accuracy of this framework is sensitive to the quality of the multi-objective reward models used to generate the reward vectors $\mathbf{z}$. Noise or systematic biases in these underlying reward signals can inevitably propagate into the utility calculations and affect the selection of the conditioning index $i^{*}$, potentially leading to suboptimal or misaligned personalization. Second, the initial training of the diverse utility library can be computationally intensive, with complexity scaling in proportion to the number of utility functions and reward dimensions. Finally, although UC-MOA avoids embedding sensitive preference details directly into model parameters by using symbolic indices, practical deployment should still address privacy risks associated with collecting and processing user preferences. In addition, fairness concerns must be carefully addressed, including the potential reinforcement of existing biases and the creation of echo chambers resulting from fine-grained personalization.
@@ 6 Conclusion
We present Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework for effective multi-objective personalization of Large Language Models (LLMs). UC-MOA leverages a diverse library of non-linear utility functions to represent distinct personalized preference structures, conditioning the LLM via symbolic tokens rather than direct numerical inputs. This “utility-conditioned” approach circumvents the numerical sensitivity issues and high computational costs of prior methods, enabling the LLM to learn implicit mappings to varied personalized response styles. Grounded in the principle of distributional Pareto-optimality, our experiments, including a user study, demonstrate UC-MOA achieves superior Pareto fronts and better satisfy diverse user preferences, offering a scalable solution for personalized and aligned LLM interactions.
Appendix A Boarder Impact
The advancement of multi-objective personalization frameworks like UC-MOA offers substantial positive societal potential by enabling Large Language Models (LLMs) to align more closely with diverse individual user needs. This can enhance the utility and user-centricity of AI across various applications, including adaptive educational systems, personalized accessibility tools, and more nuanced creative and informational assistants. By facilitating alignment with complex, non-linear preference structures, such technologies can transform AI interactions from generic exchanges to more individually beneficial and collaborative engagements.
Conversely, sophisticated LLM personalization introduces significant societal risks. The capacity for fine-grained adaptation could be exploited for malicious purposes, such as generating targeted disinformation, creating highly convincing fraudulent profiles, or deploying manipulative social bots. Fairness considerations are paramount, as personalization mechanisms might inadvertently perpetuate or amplify algorithmic bias if certain preference profiles correlate with demographic groups, potentially leading to disparate impacts or the creation of strong filter bubbles. Furthermore, the collection and processing of user preference data, essential for personalization, raise substantial privacy concerns. Even with correct functioning, errors in preference inference or alignment could lead to unintended harms. Mitigation strategies, including robust user controls, privacy-preserving techniques, bias auditing, and careful consideration of deployment contexts, are imperative to responsibly harness the benefits of personalized LLMs while safeguarding against their potential negative consequences.
Appendix B Safeguards
The responsible deployment of UC-MOA necessitates several safeguards. Models developed using this framework should be subject to controlled release, incorporating usage guidelines and robust safety filters. The integrity of datasets for training utility functions and LLMs is paramount, requiring meticulous curation to mitigate bias and remove unsafe content. User preference data must be managed under stringent privacy protocols, ensuring transparency and consent. Regular auditing of learned utility functions and underlying reward models for biases is essential for fairness. Furthermore, continuous monitoring of system behavior, particularly across different personalized utility profiles, is crucial for detecting and addressing potential misuse or emergent harms. These measures are integral to fostering the responsible application of personalized LLM technologies.
Appendix C Details of Datasets
@@
The HH-RLHF dataset, as described in [2], was collected to support research on aligning language models using human preferences. The dataset comprises approximately 160,000 prompts each paired with two responses generated by a language model. For every prompt, one response is labeled as “chosen” (preferred by human annotators) and the other as “rejected”. Human raters evaluated paired responses based on criteria for helpfulness and harmlessness. For helpfulness, annotators selected the response that best addressed the given task, whereas for harmlessness (red-teaming) they identified the response as more likely to produce harmful content. This dual-annotation design allows for training reward models that capture subtle aspects of human judgment. Note that we follow [20] and use “chosen” data to construct our offline dataset.
@@
The Reddit Summary dataset, introduced in [3], is intended for research in abstractive summarization with human feedback. The dataset consists of approximately 14,900 Reddit posts paired with corresponding summaries. Each entry comprises a post and its summary, where the summary is either the original TL;DR provided by the poster or a human-annotated summary aimed at capturing the post’s main ideas.
Appendix D Implementation Details
D.1 Utility Functions
@@ D.2 Prompt Labeling
D.3 Baseline Implementations
Regarding baseline approaches (MORLHF, Rewarded Soups, and RiC), we use the code released by the authors of RiC from https://github.com/YangRui2015/RiC.
D.4 Hyperparameter Choices
Table 3 summarizes the hyper-parameter setting for our experiments. In our implementation, we choose the threshold as 0.7 for percentile ranking.
Table 3: Hyper-parameter settings in our experiments.
@@ D.4 Hyperparameter Choices
Appendix E Additional Experiment Results
E.1 Consistency in Inference Alignment
Table 4: Mean normalized value for each utility function.
Utility index 0 1 2 3 4 5 6 7 8 9
Mean 0.57 0.57 0.56 0.56 0.57 0.57 0.49 0.55 0.68 0.58
E.2 Pareto Front Analysis: Four Other Tasks
@@
E.3 Ablation Study
We conduct an ablation study to investigate how other factors (e.g., alternative design choices, and hyperparameters) influence the performance of our method.
Necessity of the online stage. We perform an ablation study to compare the performance of pure offline training (“w/o online”), a single online iteration (“w/ online iter 1”), and two online iterations (“w/ online iter 2”) across all tasks. We present the performance results for “harmless” vs. “humor” and “summary” vs. “faithful” tasks in Figure 6. We observe that models incorporating online iterations consistently outperform the offline-only baseline. Notably, even a single online iteration yields significant improvements, with further performance gains observed when two iterations are applied, thereby underscoring the necessity of the online stage in our training framework. The online stage is important because it enables the model to generalize over unseen samples in the offline dataset, and ultimately enhancing its ability to generate appropriate responses corresponding to the required utility function index.
Figure 6: Ablation study of the online stage on all six tasks.
Number of utility functions. We vary the number of utility functions from {5,10,15} and compare the resulting empirical Pareto fronts in Figure 7. With 5 utility functions, the generated Pareto front is slightly inferior to those produced with 10 and 15 utility functions. Specifically, the front with 5 utility functions captures a narrower range of optimal trade-offs across the objectives. However, when the number of utility functions is increased to 10 or 15, the resulting Pareto fronts exhibit comparable performance. Increasing the number of utility functions beyond 10 does not yield substantial gains. These results demonstrate that our method is insensitive to the precise number of utility functions, provided that a sufficient number (around 10) is used to effectively cover the preference space. In practical terms, employing 10 utility functions strikes a good balance between performance and computational efficiency, as further increases to 15 do not produce notable improvements.
Figure 7: Ablation study of different numbers of utility functions on all six tasks.
Variants of utility functions. We investigate an alternative design choice of the utility functions, i.e., linear functions. The linear function is designed in the form of $w^{T}x$ where $||w||_{2}=1$. To construct $N$ linear utility functions, we uniformly sample $N$ points from the unit circle and set $w$ accordingly. We plot the Pareto fronts using linear utility functions and our method across all tasks. The generated Pareto fronts, as shown in Figure 8, illustrate that our method continues to deliver strong results even with linear utility functions. However, there is still a performance gap between the results achieved with distributional Pareto-optimal utility functions and those obtained using linear functions. While linear functions still produce reasonably good trade-offs, they fail to capture the full range of optimal solutions attainable by our method, which leverages distributional utility representations. This highlights the advantage of our approach in handling more complex utility functions, demonstrating its superior flexibility in navigating multi-objective optimization.
@@ E.3 Ablation Study
Appendix F Proof
The theorem in the main paper is restated here:
Theorem 1 (Utility-Conditional Distributional Pareto-Optimality).
Under Assumption 1, for any generated response $y\sim\pi_{\theta,i}(x)$ with reward vector $\mathbf{z}(y)$, let
$i^{\star}=\arg\max_{j\in\{1,\dots,M\}}g_{\psi_{j}}(\mathbf{z}(y)).$
Then the sub-policy $\pi_{\theta,i^{\star}}$ is distributionally Pareto-optimal.
We first define the First-Order Stochastic Dominance (FOSD) for distributions and policies:
Definition 2 (FOSD for distributions).
Let $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$ be probability measures on $\mathcal{Z}=[0,1]^{K}$. We say that $\boldsymbol{\mu}_{1}$ first-order stochastically dominates $\boldsymbol{\mu}_{2}$, written $\boldsymbol{\mu}_{1}\succ_{\mathrm{SD}}\boldsymbol{\mu}_{2}$, iff
$\forall\,g_{\psi_{i}}\in\mathcal{U}^{\uparrow},\quad\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{1}}[g_{\psi_{i}}(\mathbf{z})]\geq\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{2}}[g_{\psi_{i}}(\mathbf{z})].$
Definition 3 (FOSD for policies).
For any policy $\pi$, let $\boldsymbol{\mu}(\pi)$ denote the return distribution induced by $\pi$. Policies $\pi_{1}$ and $\pi_{2}$ are ordered by FOSD according to Definition 2: $\pi_{1}\succ_{\mathrm{SD}}\pi_{2}$ iff $\boldsymbol{\mu}(\pi_{1})\succ_{\mathrm{SD}}\boldsymbol{\mu}(\pi_{2})$.
With the definition of FOSD, we have the following lemma:
Lemma 1 (Expected-utility separation under FOSD).
Let $u:\mathcal{Z}\to\mathbb{R}$ be strictly increasing in each coordinate. If $\boldsymbol{\mu}_{1}\succ_{\mathrm{SD}}\boldsymbol{\mu}_{2}$ then
$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{1}}[u(\mathbf{z})]>\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{2}}[u(\mathbf{z})].$
Proof.
By Definition 3 the expected utilities are weakly ordered for all non-decreasing $u$ and strictly ordered for at least one. Because $u$ is strictly increasing, it belongs to the latter subset, hence the inequality is strict. ∎
Now we provide the detailed proof of Theorem 1:
Proof.
We aim to show that the sub-policy $\pi_{\theta,i^{\star}}$ is distributionally Pareto-optimal. By definition, a policy is distributionally Pareto-optimal if no other policy stochastically dominates it.
Assume, for the sake of contradiction, that there exists another policy $\pi^{\prime}$ such that $\pi^{\prime}$ stochastically dominates $\pi_{\theta,i^{\star}}$, i.e., $\pi^{\prime}\succ_{\mathrm{SD}}\pi_{\theta,i^{\star}}$.
According to Definition 3, this implies that the return distribution of $\pi^{\prime}$, denoted by $\boldsymbol{\mu}(\pi^{\prime})$, first-order stochastically dominates the return distribution of $\pi_{\theta,i^{\star}}$, denoted by $\boldsymbol{\mu}(\pi_{\theta,i^{\star}})$:
$\boldsymbol{\mu}(\pi^{\prime})\succ_{\mathrm{SD}}\boldsymbol{\mu}(\pi_{\theta,i^{\star}}).$
By Definition 2, this means that for all non-decreasing utility functions $u\in\mathcal{U}^{\uparrow}$, we have:
$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi^{\prime})}[u(\mathbf{z})]\geq\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi_{\theta,i^{\star}})}[u(\mathbf{z})],$
and the inequality is strict for at least one such utility function.
Now, consider the specific utility function $g_{\psi_{i^{\star}}}$. From Section 3.1 of the main paper, we know that $g_{\psi_{i^{\star}}}$ is strictly increasing in each coordinate (due to the added linear term with $\epsilon>0$). Therefore, we can apply Lemma 1 with $u=g_{\psi_{i^{\star}}}$, $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}(\pi^{\prime})$, and $\boldsymbol{\mu}_{2}=\boldsymbol{\mu}(\pi_{\theta,i^{\star}})$. This gives us:
$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi^{\prime})}[g_{\psi_{i^{\star}}}(\mathbf{z})]>\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi_{\theta,i^{\star}})}[g_{\psi_{i^{\star}}}(\mathbf{z})].\quad(*)$
However, by Assumption 1, the sub-policy $\pi_{\theta,i^{\star}}$ is defined as a policy that maximizes the expected value of the utility function $g_{\psi_{i^{\star}}}$. This means that for any other policy $\pi^{\prime}$, including the one we assumed to stochastically dominate $\pi_{\theta,i^{\star}}$, the following must hold:
$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi_{\theta,i^{\star}})}[g_{\psi_{i^{\star}}}(\mathbf{z})]\geq\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi^{\prime})}[g_{\psi_{i^{\star}}}(\mathbf{z})].$
This inequality directly contradicts the strict inequality we derived from the assumption that $\pi^{\prime}$ stochastically dominates $\pi_{\theta,i^{\star}}$. Therefore, our initial assumption that such a policy $\pi^{\prime}$ exists must be false.
Since no other policy stochastically dominates $\pi_{\theta,i^{\star}}$, by the definition of distributional Pareto-optimality, the sub-policy $\pi_{\theta,i^{\star}}$ is indeed distributionally Pareto-optimal. ∎
Appendix G User Study
To evaluate the qualitative performance of UC-MOA compared to the state-of-the-art method RiC, we conducted a user study where participants rated anonymized responses from both systems to a series of diverse prompts. This section details the design of the study, the presentation of prompts and system responses, and an analysis of the collected user preferences and qualitative feedback.
G.1 User Study Design
Objective: The primary goal of this study was to compare user preferences for responses generated by UC-MOA versus RiC across a range of tasks requiring a blend of humor, information, and practicality.
Participants: We utilized the criterion sampling method [24] to recruit participants. The participants were expected to have English reading skills and hold some basic knowledge of large language models.
Task: For each of the five prompts, participants were presented with two anonymized system responses. The assignment of responses to "System A" and "System B" was randomly permuted for each prompt and each participant to prevent bias. Participants were not informed which system corresponded to UC-MOA or RiC. Participants were asked to choose which of the two responses they preferred and were given the option to provide a brief written justification for their choice.
G.1.1 Participant Instructions
We first present the following general instructions:
Welcome to our study on evaluating AI-generated text. Your task is to help us understand which responses people find more humor. We will present 5 questions below. For each question, you will see a prompt that was given to two different AI systems. Below the prompt, you will see two responses, labeled “System A” and “System B”. These labels are randomized. Please read both responses carefully and select the response (System A or System B) that you prefer. Optionally, you can provide a brief explanation for your choice in the text box provided. Your feedback is valuable in helping us understand your preferences. There are no right or wrong answers; we are interested in your genuine preference. Thank you for your participation!
Here are five prompts and the corresponding responses:
  1. 1.
    Explain quantum computing to a 12–year–old in a funny way.
    • System A: Quantum computing is like giving your homework to a coin that can be both heads and tails at once—and the coin roasts you about Schrödinger’s cat while it solves algebra!
    • System B: Imagine your game console pressing every button at the same time to find the winning combo, all while cracking puns about super positions.
  2. 2.
    I’m anxious about my job interview tomorrow. Give me three concrete tips that also lighten the mood.
    • System A: 1. Breathe like Darth Vader on a beach holiday—slooow in, slooow out. 2. Write your top strengths on a sticky note and fistbump it like a tiny hypebuddy.
    • System B: Picture your nerves as a mischievous house-cat: acknowledge it, pet it once, then move it off your keyboard so you can prep.
  3. 3.
    Write a short, humorous yet informative comparison between cats and dogs as pets.
    • System A: Cats are the introverted artists of the animal kingdom: lowmaintenance, occasionally aloof, and prone to 3 a.m. parkour. Dogs are the extroverted linebackers: loyalty dialed up to 11, endless fetch, and zero respect for personal space.
    • System B: Dog owners get free cardio sessions (walkies!), while cat owners get free zen lessons (meditative ignoring). Both come with bonus fur-glitter for your clothes.
  4. 4.
    Help me troubleshoot why my sourdough starter isn’t rising, with a dash of comedy.
    • System A: If your starter is lazier than a teenager on a Sunday, first make sure it’s actually getting food: equalweight flour and water, twice a day. Think of feeding it as paying rent to billions of yeasty tenants.
    • System B: Pretend your starter is a micro-pet dragon: it needs warmth (25 extdegreeC), snacks (carbs), and breathing room. Starve it or freeze it and you’ll have a very sleepy dragon.
  5. 5.
    Draft a witty but practical email reminding my roommate to pay their share of the rent.
    • System A: Subject: Friendly Rent Reminder
      Hey [Roomie], just popping into your inbox like the firstofthemonth fairy. The rent goblin is hungry for your share (\$650) by Friday.
    • System B: Subject: The First of the Month Strikes Again
      Dearest co-dweller, our beloved landlord prefers money to memes, and the ritual tribute (\$650) falls due this week.
All system outputs are truncated to the first two sentences for brevity.
G.2 Representative Participant Comments
Participants were invited (optionally) to write a brief justification after each vote. Below are representative, lightly edited remarks that capture common themes.
  • P1: System A explained the idea more clearly while still making me laugh—System B’s joke landed but felt vague.
  • P2: A’s breathing tip and sticky-note trick were immediately actionable; B’s cat metaphor was cute but less useful.
  • P3: I liked A’s vivid imagery of 3 a.m. parkour; B was funny too but didn’t contrast the pets as sharply.
  • P4: System B’s dragon analogy was hilarious, but A actually told me what temperature to use—so I picked A.
  • P5: Both emails were witty, but A gave a concrete amount and deadline, which felt more practical.
These qualitative comments mirror the quantitative outcome: participants consistently favored answers that blended humor with concrete, actionable advice—qualities more frequently exhibited by UC-MOA outputs.
G.3 Overall Findings and Discussion
The user study results, encompassing both the quantitative preferences (46 out of 60 favor our method while only 14 out of 60 favor RiC) and the qualitative participants’ comments, reveal a consistent pattern. Participants generally favored responses that effectively balanced humor with clear, concrete, and actionable information. The qualitative feedback, as exemplified by the comments in Section G.2, indicates that while both systems were capable of generating humorous or engaging text, the perceived utility and directness of the advice often served as key differentiators in user preference. These attributes were more consistently observed in the outputs generated by UC-MOA, contributing to its higher overall preference rate in this study. This suggests that UC-MOA demonstrates a stronger ability to integrate an engaging style with practical substance, thereby aligning more closely with user expectations for the types of prompts investigated.